# Lower Semicontinuity of Restricted Holonomy Groups

### v1

@@ Abstract.
A result on the monotony of holonomy groups of metric connections in limits is proven. The main outcome is that given a sequence of metric connections that are Lipschitz continuous and converging in $C^{0}$, if the connections in the sequence all have holonomy contained in a closed group $H$, then also the limit connection has holonomy contained in $H$. In particular, this finds an application for Riemannian holonomy groups, where it is proven that the map assigning to a Riemannian metric the conjugacy class of its restricted holonomy group is lower semicontinuous with respect to the $C^{1}$-topology on the space of metrics.
@@ 1. Introduction
An important notion in differential geometry is the one of holonomy. Given any connection on a principal or vector bundle and a basepoint, one can define the holonomy group of the connection as the group of endomorphisms of the fibres over this point given by parallel translation along piecewise-differentiable loops. Of particular interest are the holonomy groups of the Levi-Civita connection on a Riemannian manifold, which are subgroups of the special orthogonal group. In many cases, a “special” holonomy of a Riemannian manifold implies interesting consequences for the resulting metric. For example, a smooth $4m$-dimensional Riemannian manifold with holonomy contained in the group $\mathrm{Sp}(m)\mathrm{Sp}(1)$ is automatically an Einstein manifold (see e.g. [3], Theorem 14.39).
@@
In this work, we study the changes of the holonomy group under changes of the connection. We focus on the case of metric connections on a vector bundle and prove a result of monotony of the holonomy group in limits, Theorem 2.4 below. Roughly speaking, it is shown that in the case of $C^{0}$-convergence of metric connections $\nabla^{k}$ to a metric connection $\nabla$, the connection $\nabla$ can only have a more special holonomy than the connections $\nabla^{k}$, but not a more generic one.
In particular, this finds an application for Riemannian holonomy groups. In this case, we also obtain a prettier (but slightly weaker) reformulation of the statement in terms of the lower semicontinuity of the map assigning to a $C^{1}$-Riemannian metric the conjugacy class of its restricted holonomy group.
@@
The motivation and some of the ideas of this paper are taken from the recent preprint [6] by Müller. However, the proof of the main theorem is a different one than in the cited preprint, and the resulting theorem is a stronger one (note that in [6], $C^{2}$-convergence is required instead of only $C^{1}$-convergence).
2. Monotony of the holonomy group in limits
@@
Let $M$ be a connected manifold of class $C^{1}$, let $x\in M$ and let $E\to M$ be an orientable vector bundle over $M$ with standard fibre $\mathbb{R}^{l}$, $l\in\mathbb{N}$. Let $\nabla$ be a connection on $E$.
@@
    We denote by $\Omega_{\mathrm{p}-C^{1}}(x)$ the set of all piecewise-differentiable loops
@@
    Let $c\in\Omega_{\mathrm{p}-C^{1}}(x)$. We define the parallel transport $P_{c}^{\nabla}$ along $c$ (with respect to the connection $\nabla$ on $E$) as follows: Given any $e\in E_{x}$, there exists a unique section $s$ of $E|_{c([0,1])}$ with $\nabla_{\dot{c}(t)}s(t)=0$ for all $t\in[0,1]$, and we let $P_{c}^{\nabla}(e):=s(1)$. Then
    $\mathrm{Hol}_{x}(\nabla):=\left\{P_{c}^{\nabla}\,\middle|\,c\in\Omega_{\mathrm{p}-C^{1}}(x)\right\}$
@@
    We define the restricted holonomy group of $\nabla$ at $x$,
    $\mathrm{Hol}_{x}^{0}(\nabla):=$ $\,\left\{P_{c}^{\nabla}\,\middle|\,c\in\Omega_{\mathrm{p}-C^{1}}(x)\textnormal{ such that $c$ is contractible}\right\}$
@@
    Both $\mathrm{Hol}_{x}(\nabla)$ and $\mathrm{Hol}_{x}^{0}(\nabla)$ are Lie subgroups of $\mathrm{GL}(E_{x})$. It can be shown (see e.g. [1], Section 2, Lemma 4) that $\mathrm{Hol}_{x}^{0}(\nabla)$ is the connected component of the identity in $\mathrm{Hol}_{x}(\nabla)$.
@@
    Choosing any basis of $E_{x}$, we may identify $E_{x}$ with $\mathbb{R}^{l}$ and consequently view $\mathrm{Hol}_{x}(\nabla)$ and $\mathrm{Hol}_{x}^{0}(\nabla)$ as subgroups of $\mathrm{GL}(l)$. However, this is only up to conjugation; the choice of different bases yields subgroups conjugate in $\mathrm{GL}(l)$.
Note that since any $C^{1}$-map is in particular Lipschitz, a manifold of class $C^{1}$ is a Lipschitz manifold, that is, the transition maps of its defining atlas are bi-Lipschitz ([8], Definition 1). It is well-defined to speak of locally Lipschitz continuous functions on such a manifold.
In the present work, we will be concerned with the following setting.
Situation 2.2.
  • Let $M$ be a connected manifold of class $C^{1}$, let $x\in M$ and $E\to M$ be an orientable vector bundle with standard fibre $\mathbb{R}^{l}$. We denote by $E^{*}$ the dual bundle of $E$ and by $\mathrm{S}^{2}E^{*}$ the symmetrized tensor product of the dual bundle.
  • Let $g\in C^{1}(M,\mathrm{S}^{2}E^{*})$ be a bundle metric on $E$ (that remains fixed throughout this section unless otherwise stated). Let $\nabla^{g}$ be a connection on $E$ whose coefficients are locally Lipschitz continuous and that is compatible with $g$, i. e. $X(g(v,w))=g(\nabla^{g}_{X}v,w)+g(v,\nabla^{g}_{X}w)$ for all $X\in TM$, $v,w\in C^{1}(M,E)$.
  • Given any metric $h$ on $E$, let
@@
    be the special orthogonal group of $E_{x}$ with respect to the metric $h|_{x}$. Then, since $\nabla^{g}$ is metric, we have $P_{c}^{\nabla^{g}}\in\mathrm{SO}_{l}(E_{x})_{g}$ for all $c\in\Omega_{\mathrm{p}-C^{1}}(x)$.
  • An orthonormal basis of $E_{x}$ with respect to some metric $h$ are vectors $e_{1},\dots,e_{l}$ with
    $\forall i,j=1,\dots,n:h(e_{i},e_{j})=\delta_{ij},$
    where $\delta_{ij}$ is the Kronecker delta. We choose a positively oriented orthonormal basis $(b_{1},\dots,b_{l})$ of $E_{x}$ with respect to the metric $g$ and use it to identify $E_{x}$ with $\mathbb{R}^{l}$. This then identifies $\mathrm{SO}_{l}(E_{x})_{g}$ with $\mathrm{SO}(l)\subset\mathrm{GL}(l)$ and consequently we may view $\mathrm{Hol}_{x}(\nabla^{g})$ as an explicit subgroup of $\mathrm{SO}(l)$.
@@
    Using the basis $(b_{1},\dots,b_{l})$ of $E_{x}$, we may view all bilinear forms $Q\in\mathrm{S}^{2}E^{*}_{x}$ and all endomorphisms $Z\in\mathrm{End}(E_{x})$ as $l\times l$-matrices over $\mathbb{R}$. In particular, we may associate to a bilinear form $Q$ its representing matrix $M_{Q}$ and via this an endomorphism $Z_{Q}$ of $E_{x}\cong\mathbb{R}^{l}$. [^1] To be very explicit, therefore $Q(X,Y)=g(X,Z_{Q}Y)$ for all $X,Y\in E_{x}$, and if we view $X,Y$ also as their respective coordinate vectors with respect to the basis $(b_{1},\dots,b_{l})$, then this is nothing but $X^{T}M_{Q}Y$. This defines an embedding $\mathrm{S}^{2}E^{*}_{x}\to\mathrm{End}(E_{x})$, $Q\mapsto Z_{Q}$. Now $\mathrm{End}(E_{x})$ can be turned into a normed vector space by using the operator norm induced by the metric $g$. By pullback along the embedding described above, also $\mathrm{S}^{2}E^{*}_{x}$ becomes a normed vector space. Unless otherwise stated, in the following these norms will be used. (As all vector spaces involved are finite-dimensional, the resulting topologies do not depend on $g$ or the basis.)
  2. (2)
    Let $h\in C^{1}(M,\mathrm{S}^{2}E^{*})$ with $h|_{x}\in B_{1}(g|_{x})$. Let $M_{h}$ be the representing matrix of $h|_{x}$ and $Z_{h}$ the corresponding endomorphism as in the previous point. We define
    $\mathrm{SO}(l)_{h}:=\{A\in\mathrm{GL}(l)\,|\,A^{T}M_{h}A-M_{h}=0\}$
    (the $x$-dependency is understood). Let $W_{h}$ be a symmetric positive definite square root of $Z_{h}$ as defined by the power series of $\sqrt{1+c}$ in $B_{1}(0)$. Then, the map $\mathrm{SO}(l)\to\mathrm{SO}(l)_{h},A\mapsto W_{h}^{-1}AW_{h}$ is an isomorphism. [^2] Here we again use the basis $(b_{1},\dots,b_{l})$ of $E_{x}$ to identify $E_{x}$ with $\mathbb{R}^{l}$, $\mathrm{SO}(l)$ with a matrix group, $Z_{h}$ with $M_{h}$ and $W_{h}$ also with an explicit $l\times l$-matrix over $\mathbb{R}$. Indeed this follows from the following matrix computation: If $A\in\mathrm{SO}(l)$, then
    $(W_{h}^{-1}AW_{h})^{T}h|_{x}$ $(W_{h}^{-1}AW_{h})-M_{h}$
    $=W_{h}^{T}A^{T}W_{h}^{-T}W_{h}^{T}W_{h}W_{h}^{-1}AW_{h}-M_{h}$
    $=W_{h}^{T}A^{T}AW_{h}-M_{h}$
    $=W_{h}^{T}W_{h}-M_{h}=M_{h}-M_{h}=0,$
@@
    thus $W_{h}^{-1}AW_{h}\in\mathrm{SO}(l)_{h}$; the inverse isomorphism is simply given by $A\mapsto W_{h}AW_{h}^{-1}$. We call this isomorphism the nonstandard embedding $\phi_{h}:\mathrm{SO}(l)\to\mathrm{SO}(l)_{h}\subset\mathrm{GL}(l)$. By the description of $W_{h}$ in terms of a power series in $h$, we obtain that if we have a sequence $(h_{k})_{k\in\mathbb{N}}\subset C^{1}(M,\mathrm{S}^{2}E^{*})$ with $h_{k}|_{x}\in B_{1}(g|_{x})\cap\mathrm{S}^{2}E_{x}^{*}$ for all $k$ and $h_{k}|_{x}\to g|_{x}$ in $\mathrm{S}^{2}E^{*}_{x}$, then $W_{h_{k}}\to I_{l}$ (the $l\times l$-identity matrix).
  3. (3)
    Given any two connections $\nabla,\tilde{}\nabla$ on $E$ with continuous coefficients, its difference $\nabla-\tilde{}\nabla$ is a continuous section of $T^{*}M\otimes E\otimes E^{*}$. We equip the space $\Gamma(T^{*}M\otimes E\otimes E^{*})$ of such sections with the compact-open $C^{r}$-topology, $r=0$, which is defined as follows ([4], Chapter 2.1): Given any two charts $\mathbf{x}:U\to V,\mathbf{v}:U^{\prime}\to V^{\prime}$ of $M$ respectively $T^{*}M\otimes E\otimes E^{*}$, a compact $K\subseteq U$, $\epsilon>0$ and $f\in\Gamma(T^{*}M\otimes E\otimes E^{*})$, we let
    $B(f,\mathbf{x},\mathbf{v},K,\epsilon):=\Big\{$ $g\in\Gamma(T^{*}M\otimes E\otimes E^{*})\,|\,g(K)\subseteq V,\forall 0\leq k\leq r:$
    $\left\lVert D^{k}(\mathbf{v}\circ f\circ\mathbf{x}^{-1})-D^{k}(\mathbf{v}\circ g\circ\mathbf{x}^{-1})\right\rVert_{C^{0}}<\epsilon\Big\}.\qquad\qquad$
    The set of all such $B(f,\mathbf{x},\mathbf{v},K,\epsilon)$ has the basis property and thus defines a topology on $\Gamma(T^{*}M\otimes E\otimes E^{*})$. Analogously, we define compact-open $C^{r}$-topologies, $r\geq 0$, on the spaces of sufficiently often differentiable sections of other bundles (if $M$ is of class at least $C^{r}$).
    Using this, we can speak of $C^{0}$-convergence in $\Gamma(T^{*}M\otimes E\otimes E^{*})$, and we say that a sequence of connections $\nabla^{k}$ converges in $C^{0}$ to a connection $\nabla$ on $E$ if their difference converges to zero. If this is the case, then the restrictions of $\nabla^{k}$-$\nabla$ to any compact $K\subseteq M$ converge to zero uniformly on $K$ (with respect to any set of charts covering $K$).
Now, the announced result can be stated.
Theorem 2.4.
Assume Situation 2.2 and let $H\subseteq\mathrm{SO}(l)$ be closed. Moreover, for $k\in\mathbb{N}$, let $g_{k}\in C^{1}(M,\mathrm{S}^{2}E^{*})$ be a bundle metric on $E$ and $\nabla^{k}$ be a connection with locally Lipschitz continuous coefficients compatible with $g_{k}$, such that $g_{k}|_{x}\to g|_{x}$ in $\mathrm{S}^{2}E^{*}_{x}$ and $\nabla^{k}-\nabla^{g}\to 0$ in the compact-open $C^{0}$-topology on $\Gamma(M,T^{*}M\otimes E^{*}\otimes E)$.
Assume that for all $k\in\mathbb{N}$, there exists a positively oriented $g_{k}$-orthonormal basis of $E_{x}$ such that if we use this basis to identify $E_{x}$ with $\mathbb{R}^{l}$, we have $\mathrm{Hol}_{x}(\nabla^{k})\subseteq H$. Then, there exists a positively oriented $g$-orthonormal basis $(c_{1},\dots,c_{l})$ of $E_{x}$ such that if we use this basis to identify $E_{x}$ with $\mathbb{R}^{l}$, we have $\mathrm{Hol}_{x}(\nabla^{g})\subseteq H$.
@@
Note first that the conclusion of the theorem may be reformulated as follows: If we use our pre-chosen basis $(b_{1},\dots,b_{l})$ to identify $E_{x}$ with $\mathbb{R}^{l}$, then there exists $U\in\mathrm{SO}(l)$ such that $U\mathrm{Hol}_{x}(\nabla^{g})U^{-1}\subseteq H$. [^3] Indeed: It suffices to take $U$ as the matrix such that $\sum_{k=1}^{l}U_{ki}c_{k}=b_{i}$ for $i=1,\dots,l$. Therefore, we fix once and for all the basis $(b_{1},\dots,b_{l})$ and use it to identify $E_{x}$ with $\mathbb{R}^{l}$.
We may assume that $g_{k}|_{x}\in B_{1}(g|_{x})\cap\mathrm{S}^{2}E_{x}^{*}$ for all $k\in\mathbb{N}$ by dropping a finite number of elements from the sequence, hence we may define the matrices $W_{g_{k}}$, the subgroups $\mathrm{SO}(l)_{g_{k}}$ and the nonstandard embedding $\phi_{g_{k}}$ for $k\in\mathbb{N}$ as in Definition 2.3. Furthermore, we denote by $H_{g_{k}}$ the image of the composition of the inclusion $H\hookrightarrow\mathrm{SO}(l)$ and the nonstandard embedding $\mathrm{SO}(l)\to\mathrm{SO}(l)_{g_{k}}$.
The assumption then means that for all $k\in\mathbb{N}$ there exists an $U_{k}\in\mathrm{SO}(l)_{g_{k}}$ such that $U_{k}\mathrm{Hol}_{x}(\nabla^{k})U_{k}^{-1}\subseteq H_{g_{k}}$. [^4] Note that since the connection $\nabla^{k}$ is metric with respect to $g_{k}$, we have $\mathrm{Hol}_{x}(\nabla^{k})\subseteq\mathrm{SO}(l)_{g_{k}}$ for all $k\in\mathbb{N}$. Now, the sequence $\phi_{g_{k}}^{-1}(U_{k})$ is contained in the compact group $\mathrm{SO}(l)$ and hence possesses a convergent subsequence. Moreover, using the explicit description of $\phi_{g_{k}}$ in terms of conjugation by matrices converging to the identity matrix, we see that if $(\phi_{g_{k_{m}}}^{-1}(U_{k_{m}}))_{m\in\mathbb{N}}$ converges to some $U\in\mathrm{SO}(l)$, then also $(U_{k_{m}})_{m\in\mathbb{N}}$ converges to $U$.
Therefore, passing to the subsequence without changing the notation for brevity, we may assume that $(U_{k})_{k\in\mathbb{N}}$ converges to some $U\in\mathrm{SO}(l)$. At this point, it suffices to prove the following claim:
@@
Proof of the claim. Let $c\in\Omega_{\mathrm{p}-C^{1}}(x)$. By definition of the compact-open topology, the restrictions of the connections to the (compact) image of $c$ converge uniformly in $\Gamma((T^{*}M\otimes E^{*}\otimes E)|_{c([0,1])})$. Therefore, the coefficients of the parallel transport equation defining $P_{c}^{\nabla^{k}}$ converge uniformly to the coefficients of the parallel transport equation defining $P_{c}^{\nabla^{g}}$. As the solutions to ordinary differential equations depend continuously on their coefficients, we have $\lim\limits_{k\to\infty}P_{c}^{\nabla^{k}}=P_{c}^{\nabla^{g}}$. [^5] See e. g. [7] Satz 4.1.2; to be precise, the “continuous dependence on the coefficients” is the statement that if the coefficients converge in $C^{0}$ (and are all locally Lipschitz continuous), then the solutions converge in $C^{0}$. Note that we can assume that the coefficients of the parallel transport equations along $c$ are continuous since we can parametrize $c$ such that it has vanishing speed at the points where its parametrization by arc-length would not be differentiable. Moreover, then the pullback of the locally Lipschitz continuous coefficients of $\nabla^{k}$ along the differentiable path $c$ is again locally Lipschitz continuous. Thus, also

### v2

# Monotonicity of Holonomy Groups

@@ Abstract.
We prove the following monotonicity result for the holonomy group: Given a sequence of metric connections converging in $C^{0}$ such that all its members have holonomy contained in a closed group $H$, also their limit connection needs to have holonomy contained in $H$. As a corollary, for a sequence of Riemannian metrics converging in $C^{1}$ and having special restricted holonomy, their limit metric must also have special restricted holonomy. In particular, this implies that the map assigning to Riemannian metrics on a manifold the conjugacy classes of their restricted holonomy groups is lower semicontinuous with respect to the order relation given by inclusion of representatives.
@@ 1. Introduction
An important notion in differential geometry is the one of holonomy. Given a connection on a principal or vector bundle, one can define the holonomy group of the connection as the group of endomorphisms of the fibres given by parallel translation along piecewise-differentiable based loops. Of particular interest are the holonomy groups of the Levi-Civita connection on a Riemannian manifold, which are subgroups of the special orthogonal group. In many cases, a “special” holonomy of a Riemannian manifold implies interesting consequences for the resulting metric. For example, a smooth $4m$-dimensional Riemannian manifold with holonomy contained in the group $\mathrm{Sp}(m)\mathrm{Sp}(1)$ is automatically an Einstein manifold (see e. g. [3], Theorem 14.39).
@@
In this work, we study the dependence of the holonomy group on the connection. We focus on the case of metric connections on a vector bundle and prove the following monotonicity result:
Theorem 1.1.
Let $M$ be a connected manifold of class $C^{2}$, let $x\in M$ and $E\to M$ be an orientable real $C^{1}$-vector bundle with standard fibre $\mathbb{R}^{l}$. Let $g\in C^{1}(M,\mathrm{S}^{2}E^{*})$ be a bundle metric on $E$ and $\nabla^{g}$ be a $C^{0}$-connection on $E$ that is compatible with $g$. Moreover, for $k\in\mathbb{N}$, let $g_{k}\in C^{1}(M,\mathrm{S}^{2}E^{*})$ be a bundle metric on $E$ and $\nabla^{k}$ be a $C^{0}$-connection compatible with $g_{k}$, such that $g_{k}|_{x}\to g|_{x}$ in $\mathrm{S}^{2}E^{*}_{x}$ and $\nabla^{k}\to\nabla^{g}$ in $C^{0}$.
Let $H\subseteq\mathrm{SO}(l)$ be closed. Assume that for all $k\in\mathbb{N}$, there exists a positively oriented $g_{k}$-orthonormal basis of $E_{x}$ such that with respect to this basis, we have $\mathrm{Hol}_{x}(\nabla^{k})\subseteq H$. Then, there exists a positively oriented $g$-orthonormal basis of $E_{x}$ such that with respect to this basis, we have $\mathrm{Hol}_{x}(\nabla^{g})\subseteq H$.
Here, $\mathrm{S}^{2}E^{*}$ is the symmetric tensor product of $E^{*}$ with itself and $C^{0}$-convergence of connections is defined via the compact-open topology on $C^{0}(T^{*}M\otimes E\otimes E^{*})$, see Section 2. Thus, it is shown that in the case of $C^{0}$-convergence of metric connections, the holonomy can only become more special and not more generic.
In particular, this finds an application for Riemannian holonomy groups. Given a manifold $M$, denote by $\mathrm{Met}(M)$ be the set of all $C^{1}$-Riemannian metrics on $M$, equipped with the compact-open $C^{1}$-topology. We obtain:
Corollary 1.2.
Let $(M,g)$ be an oriented $n$-dimensional Riemannian manifold of class $C^{2}$, let $g\in\mathrm{Met}(M)$ and for $k\in\mathbb{N}$, let $g_{k}\in\mathrm{Met}(M)$. Let $H\subseteq\mathrm{SO}(n)$ be a closed subgroup and let $x\in M$. Assume that $g_{k}\to g$ in $\mathrm{Met}(M)$ and for all $k\in\mathbb{N}$, we have $\mathrm{Hol}_{x}(g_{k})\subseteq H$ up to conjugation in $\mathrm{SO}(n)$. Then, $\mathrm{Hol}_{x}(g)\subseteq H$ up to conjugation in $\mathrm{SO}(n)$.
These statements hold true for the full holonomy group and a fortiori also for the restricted one. However, for the latter, we can show more. Let $\mathrm{CSC}(\mathrm{SO}(n))$ be the set of closed subgroups of $\mathrm{SO}(n)$ modulo the conjugation equivalence relation; it will be shown in Section 3 that this is a partially ordered set with respect to inclusion of representatives. Then, with respect to this order relation, the above monotonicity result implies the following:
Theorem 1.3.
Let $M$ be an oriented $n$-dimensional manifold of class $C^{2}$ and let $x\in M$. The map
is lower semicontinuous.
An appropriate definition of lower semicontinuity of a map from a topological space into a partially ordered set is given in Section 3.
@@
The motivation and some of the ideas of this paper are taken from the recent preprint [7] by Müller. However, the proof of the main theorem is different from the cited preprint, and the resulting theorem is stronger (note that [7] requires $C^{2}$-regularity of the involved metrics).
2. Situation and proof of Theorem 1.1
@@
Let $M$ be a connected manifold of class $C^{2}$, let $x\in M$ and let $E\to M$ be an orientable real $C^{1}$-vector bundle over $M$ with standard fibre $\mathbb{R}^{l}$, $l\in\mathbb{N}$.
@@
    We denote by $\Omega_{\textnormal{p-}C^{1}}(x)$ the set of all piecewise-differentiable loops
@@
    Let $\nabla$ be a connection on $E$ and $c\in\Omega_{\textnormal{p-}C^{1}}(x)$. We define the parallel transport $P_{c}^{\nabla}$ along $c$ (with respect to $\nabla$) as follows: Given any $e\in E_{x}$, there exists a unique section $s$ of $E|_{c([0,1])}$ with $\nabla_{\dot{c}(t)}s(t)=0$ for all $t\in[0,1]$, and we let $P_{c}^{\nabla}(e):=s(1)$. Then
    $\mathrm{Hol}_{x}(\nabla):=\left\{P_{c}^{\nabla}\,\middle|\,c\in\Omega_{\textnormal{p-}C^{1}}(x)\right\}$
@@
    Given a connection $\nabla$ on $E$, we define its restricted holonomy group at $x$,
    $\mathrm{Hol}_{x}^{0}(\nabla):=$ $\,\left\{P_{c}^{\nabla}\,\middle|\,c\in\Omega_{\textnormal{p-}C^{1}}(x)\textnormal{ such that $c$ is contractible}\right\}$
@@
    Both $\mathrm{Hol}_{x}(\nabla)$ and $\mathrm{Hol}_{x}^{0}(\nabla)$ are immersed Lie subgroups of $\mathrm{GL}(E_{x})$. It can be shown (see e. g. [1], Section 2, Lemma 4) that $\mathrm{Hol}_{x}^{0}(\nabla)$ is the connected component of the identity in $\mathrm{Hol}_{x}(\nabla)$.
@@
    Given two connections $\nabla,\tilde{}\nabla$ on $E$ with continuous coefficients, their difference $\nabla-\tilde{}\nabla$ is a continuous section of $T^{*}M\otimes E\otimes E^{*}$. We equip the space $C^{0}(M,T^{*}M\otimes E\otimes E^{*})$ of such sections with the compact-open $C^{0}$-topology (see Appendix A) and say that a sequence of connections $\nabla^{k}$ on $E$ converges in $C^{0}$ to a connection $\nabla$ if $\nabla^{k}-\nabla\to 0\in C^{0}(M,T^{*}M\otimes E^{*}\otimes E)$.
  5. (5)
    Let $h$ be a metric on $E$ and $\nabla^{h}$ be a $C^{0}$-connection on $E$ that is compatible with $h$, i. e. $X(h(v,w))=h(\nabla^{h}_{X}v,w)+h(v,\nabla^{h}_{X}w)$ for all $X\in TM$, $v,w\in C^{1}(M,E)$. We denote by
@@
    the special orthogonal group of $E_{x}$ with respect to the metric $h|_{x}$. Then, since $\nabla^{h}$ is metric, we have $P_{c}^{\nabla^{h}}\in\mathrm{SO}_{l}(E_{x})_{h}$ for all $c\in\Omega_{\textnormal{p-}C^{1}}(x)$.
  6. (6)
    Choosing any basis of $E_{x}$, we may identify $E_{x}$ with $\mathbb{R}^{l}$ and consequently view $\mathrm{Hol}_{x}(\nabla)$ and $\mathrm{Hol}_{x}^{0}(\nabla)$ for any $\nabla$ as subgroups of $\mathrm{GL}(l)$. However, this is only up to conjugation; the choice of different bases yields subgroups conjugate in $\mathrm{GL}(l)$.
  7. (7)
    Let $g\in C^{1}(M,\mathrm{S}^{2}E^{*})$ be the bundle metric on $E$ from the statement of Theorem 1.1. We choose a $g$-orthonormal basis $(b_{1},\dots,b_{l})$ of $E_{x}$ and use it to identify $E_{x}$ with $\mathbb{R}^{l}$, $\mathrm{SO}_{l}(E_{x})_{g}$ with $\mathrm{SO}(l)$ and consequently $\mathrm{Hol}_{x}(\nabla^{g})$ with a subgroup of $\mathrm{SO}(l)$.
@@
    The basis $(b_{1},\dots,b_{l})$ allows us to associate to any bilinear form $Q$ on $E_{x}$ its representing matrix $M_{Q}\in\mathbb{R}^{l\times l}$; if $Q$ is non-degenerate, $M_{Q}$ is invertible. Given $h\in C^{1}(M,\mathrm{S}^{2}E^{*})$ with $M_{h|_{x}}\in B_{1}(M_{g|_{x}})$, we define
    $\mathrm{SO}(l)_{h}:=\{A\in\mathrm{GL}(l)\,|\,A^{T}M_{h|_{x}}A=M_{h|_{x}}\}$
@@
    (the $x$-dependency is understood). Let $W_{h}$ be a symmetric positive definite square root of $M_{h|_{x}}$ as defined by the power series of $\sqrt{1+c}$ in $B_{1}(0)$; then the map $\mathrm{SO}(l)\to\mathrm{SO}(l)_{h},A\mapsto W_{h}^{-1}AW_{h}$ is an isomorphism. We call it the nonstandard embedding $\phi_{h}:\mathrm{SO}(l)\to\mathrm{SO}(l)_{h}\subset\mathrm{GL}(l)$. For a sequence $(h_{k})_{k\in\mathbb{N}}\subset C^{1}(M,\mathrm{S}^{2}E^{*})$ with $M_{h_{k}|_{x}}\in B_{1}(M_{g|_{x}})$ for all $k$ and $M_{h_{k}|_{x}}\to M_{g|_{x}}$ in $\mathbb{R}^{l\times l}$, we have $W_{h_{k}}\to I_{l}$ (the $l\times l$-identity matrix).
Now, we can prove the first announced result, Theorem 1.1.
@@
Let $(b_{1},\dots,b_{l})$ be the positively oriented $g$-orthonormal basis of $E_{x}$; we use it to identify $E_{x}$ with $\mathbb{R}^{l}$ and $H$ with a subgroup of $\mathrm{SO}(l)$ as before. Then, the conclusion of the theorem may be reformulated as follows: There exists $U\in\mathrm{SO}(l)$ such that $U\mathrm{Hol}_{x}(\nabla^{g})U^{-1}\subseteq H$.
We may assume that the representing matrices of $g_{k}|_{x}$ and $g|_{x}$ satisfy $M_{g_{k}|_{x}}\in B_{1}(M_{g|_{x}})$ for all $k\in\mathbb{N}$ by dropping a finite number of elements from the sequence, hence we may define the matrices $W_{g_{k}}$, the subgroups $\mathrm{SO}(l)_{g_{k}}$ and the nonstandard embedding $\phi_{g_{k}}$ for $k\in\mathbb{N}$ as in Definition 2.1. Let us denote by $H_{g_{k}}$ the image of the composition of the inclusion $H\hookrightarrow\mathrm{SO}(l)$ and the nonstandard embedding $\mathrm{SO}(l)\to\mathrm{SO}(l)_{g_{k}}$.
The identification $E_{x}=\mathbb{R}^{l}$ translates the assumptions of the theorem into the assertion that for all $k\in\mathbb{N}$ there exists $U_{k}\in\mathrm{SO}(l)_{g_{k}}$ with $U_{k}\mathrm{Hol}_{x}(\nabla^{k})U_{k}^{-1}\subseteq H_{g_{k}}$. Now, the sequence $\phi_{g_{k}}^{-1}(U_{k})$ is contained in the compact group $\mathrm{SO}(l)$ and hence possesses a convergent subsequence. Moreover, using the explicit definition of $\phi_{g_{k}}$, we see that if $(\phi_{g_{k_{m}}}^{-1}(U_{k_{m}}))_{m\in\mathbb{N}}$ converges to some $U\in\mathrm{SO}(l)$, then also $(U_{k_{m}})_{m\in\mathbb{N}}$ converges to $U$.
Therefore, passing to the subsequence without changing the notation for brevity, we may assume that $(U_{k})_{k\in\mathbb{N}}$ converges to some $U\in\mathrm{SO}(l)$. It now only remains to prove the following:
@@
Proof of the claim. It is sufficient to show that for any $c\in\Omega_{\textnormal{p-}C^{1}}(x)$, we have $P_{c}^{\nabla^{g}}\in H$. By definition of the compact-open topology, the restrictions of the connections to the (compact) image of $c$ converge in $C^{0}$. Therefore, the coefficients of the parallel transport equation defining $P_{c}^{\nabla^{k}}$ converge uniformly to the coefficients of the parallel transport equation defining $P_{c}^{\nabla^{g}}$. From well-known facts about ordinary differential equations it follows that after possibly again passing to a subsequence, we may assume that $\lim\limits_{k\to\infty}P_{c}^{\nabla^{k}}=P_{c}^{\nabla^{g}}$. (See e. g. [4], Chapter I, Theorem 2.4.) Thus, also
@@
It shall be stressed that the proof of the claim above makes no assertion about the uniformity of the convergence $P_{c}^{\nabla^{k}}\to P_{c}^{\nabla^{g}}$ with respect to the loop $c$, but such one is also not required.
Remark 2.3.